Tyler's parents pay him $0.50 to do the laundry and $1.25 to mow the lawn. In one month, he does the laundry 6 more times than he mows the lawn. If his parents pay him $12.75 that month, how many times did he mow the lawn?
Answer by stanbon(75887)

Tyler's parents pay him $0.50 to do the laundry and $1.25 to mow the lawn. In one month, he does the laundry 6 more times than he mows the lawn. If his parents pay him $12.75 that month, how many times did he mow the lawn?
-------------
Let number of times he mows the lawn be "x"
# of times he does laundry = x+6
-----------------
Money from lawn work = 1.25x dollars
Money from laundry work = 0.50(x+6) = 0.5x+3 dollars
------------------
EQUATION:
money + money = 12.75 dollars
1.25x + 0.5x+3 = 12.75
1.75x = 9.75
x = 5.571...
Comment:
That doesn't sound right.
x should be a whole number.
Cheers,
Stan H.